OpenStudy (auctoratrox):

Twelve people join hands for a circle dance. Suppose 6 are men and the other 6 are women. If they alternate by gender, in how many ways can they do the dance?

OpenStudy (lgbasallote):

the men and women cannot be distinguished from one another, yes/

OpenStudy (auctoratrox):

Is there some sort of combination or permutation I can use to solve this?

OpenStudy (lgbasallote):

yes. circular permutation

OpenStudy (lgbasallote):

\[P = (n-1)!\] where n = number of entities

OpenStudy (lgbasallote):

in this case, i suggest you count one pair as one entity

OpenStudy (lgbasallote):

then count the number of pairs...then use circular permutation

OpenStudy (lgbasallote):

then multiply by 2 (because the order of the pairs are interchangeable0

OpenStudy (lgbasallote):

i suppose you did not get what i said.../

OpenStudy (auctoratrox):

I'm working it out, just hold on

OpenStudy (lgbasallote):

oh good. i thought my multiple comments confused you

OpenStudy (auctoratrox):

So I'm getting that there are 6! possible pairs of males and females

OpenStudy (lgbasallote):

there are 6 pairs.... but the circular permutation is (n-1)!

OpenStudy (lgbasallote):

in this case, your n is 6 (because you have 6 pairs, thus, 6 entities

OpenStudy (auctoratrox):

So I have 6 pairs, thus (6-1)! = 5!

OpenStudy (lgbasallote):

right. then, since the pairs are interchangeable in order, you multiply by 2!

OpenStudy (auctoratrox):

so I get 5!2!, but why am I multiplying by 2! ?

OpenStudy (lgbasallote):

because the pairs are interchangeable think of it like this for example, this is one order b g b g b g ^think of them as in a circle another order would be g b g b g b see what i mean?

OpenStudy (auctoratrox):

ok I see what you mean

OpenStudy (lgbasallote):

good.

OpenStudy (lgbasallote):

so that would be the end